Calculate the average kinetic energy and RMS velocity of a xenon atom at 166.126 K, which is just 1.075 degrees above the boiling point of xenon at 1 atmosphere pressure. The atomic mass of xenon is 131.293 amu. Because the kinetic energy of a single atom is so tiny, express the KE in units of 10−23 10-23 J.
Answer #1

Boiling point of xenon = -108.1 °C = 165.05 K
T = 165.05 + 1.075 K
T =  166.125 K

K.Eavg = 3/2 * K*T
K.Eavg = 3/2 * 1.38 * 10^-23 * 166.125
K.Eavg = 343.88 * 10^-23 J

Vrms = sqrt[(2*K.E)/m]
Vrms = sqrt[ ( 2 * 343.88 * 10^-23) / (131.293 * 1.66 * 10^-27) ]
Vrms = 177.6 m/s